The average loan of an employee is Php 18, 200. If the debt is normally distributed with the standard deviation of Php 6, 150, find the probability that the employee owes between Php 10, 000 and Php 25, 000.

1
Expert's answer
2022-03-15T05:13:39-0400

"P(10000<X<25000)=P(\frac{10000-18200}{6150}<Z<\frac{25000-18200}{6150})=P(-1.33<Z<1.11)="

"=P(Z<1.11)-P(Z<-1.33)=0.7747."